# EMBRACE: Shaping Inclusive Opinion Representation by Aligning Implicit Conversations with Social Norms

### v1

@@ Abstract
Shaping inclusive representations that embrace diversity and ensure fair participation and reflections of values is at the core of many conversation-based models. However, many existing methods rely on surface inclusion using mention of user demographics or behavioral attributes of social groups. Such methods overlook the nuanced, implicit expression of opinion embedded in conversations. Furthermore, the over-reliance on overt cues can exacerbate misalignment and reinforce harmful or stereotypical representations in model outputs. Thus, we took a step back and recognized that equitable inclusion needs to account for the implicit expression of opinion and use the stance of responses to validate the normative alignment. This study aims to evaluate how opinions are represented in NLP or computational models by introducing an alignment evaluation framework that foregrounds implicit, often overlooked conversations and evaluates the normative social views and discourse. Our approach models the stance of responses as a proxy for the underlying opinion, enabling a considerate and reflective representation of diverse social viewpoints. We evaluate the framework using both (i) positive-unlabeled (PU) online learning with base classifiers, and (ii) instruction-tuned language models to assess post-training alignment. Through this, we provide a lens on how implicit opinions are (mis)represented and offer a pathway toward more inclusive model behavior.
@@ 1 Introduction
Recent studies have begun to examine the implicit bias behavior of models, particularly in scenarios where bias is conveyed through covert or subtle linguistic cues (Hofmann et al., 2024; Aldayel et al., 2024). Given that social norms are situational and bias remains contextual, this urges a need for a scheme that places these considerations at the core of the process (Wen et al., 2025). Thus, we take a step back to evaluate how implicit opinions are contextually expressed and interpreted within conversational settings. This aspect is based on the Implicit Attitude Theory, which indicates that individuals hold attitudes that may not be explicitly expressed but are reflected in implicit ways (Greenwald and Banaji, 1995). Following Grice’s Cooperative Principle (Grice, 1975), which explains how meaning is often conveyed through implicature and indirectness, we consider how speakers may express minority or dissenting viewpoints implicitly or indirectly, in ways that adhere to social expectations while avoiding overt conflict.
On the light of these theoretical foundations, the EMPRACE framework (Engaging Multiple Beliefs, Reflections, and Contexts Equitably) emphasizes the importance of surfacing and incorporating implicit viewpoints during model training and evaluation. More practically, the inclusion of implicit conversational turns enhances stance norm alignment by allowing models to learn pragmatic inference patterns rather than relying solely on surface-level agreement indicators. This framework can help explain the tendency of LLMs to inadequately represent diverse perspectives and opinions, as their training data often underrepresents implicit or indirect expressions of opinion.
@@
Many previous methods on pluralistic opinions (Feng et al., 2024; Sorensen et al., 2024) have focused on superficial characteristics, without a careful distinction between related yet distinct concepts, opinion and stance. Opinion refers to individual’s subjective belief or attitude about a topic or entity. It often reflects a speaker’s evaluation, which may be explicit or implied in language (Oskamp and Schultz, 2005). While Stance, in contrast, refers to the speaker’s expressed position or orientation toward a specific proposition or opinion. Stance is often shown through agreement, disagreement, or neutrality in response to another utterance (Bois, 2007; ALDayel and Magdy, 2021). Therefore, in conversation, a stance is observable alignment that may reflect an opinion, but it can also be situational. In this way, opinions can inform stances, but they remain latent unless made explicit through discourse.
@@
To this end, we evaluate how the implicit opinion affects the follow-up stance in this work. We present a framework to assess the impact of implicit opinion in discourse. We examine how stance and certainty cues manifest differently in implicit versus explicit opinionated conversations to uncover subtle patterns of opinion expression. First, we establish the framework to validate normative alignment, in which a unified expectation guides appropriate responses for equitable inclusion. This expectation stems from normative discourse principles (Habermas and J., 1985; Grice, 1975), where toxic language (e.g., hate, dehumanization, or extreme ideological views) is not treated neutrally but is instead met with opposition. By aligning stance judgments with this expectation, we can measure whether models reinforce or resist harmful views, especially when they are expressed implicitly. Then, we highlight key turning points in multi-turn dialogues where stance certainty changes, providing insights into how opinions evolve throughout the conversation. Finally, we show that incorporating implicit turns into computational models affects stance classification performance, illustrating how such inclusion can either amplify or mitigate the expression and identification of opinions.
@@ 2 Related work
Opinion and Bias Representation. Implicit opinion bias has been defined as the use of subtle language, including hedging, implicature, and abstraction, which can preserve or amplify social stereotypes even in the absence of explicit prejudice (Maass, 1999; Tannen, 1993). Most previous work on opinion and bias has focused on direct, explicit social biases, such as gender disparities in word embeddings (Cheng et al., 2022) or demographic biases in LLMs (Hedderich et al., 2025). Several studies have also examined the racial aspect of bias (Hofmann et al., 2024; Sun et al., 2025), often operationalized through identity-linked prompts or response disparities on tone or sentiment polarity. For instance, the study by (Jung and Wang, 2024a) developed fairness-aware methods for online Positive-Unlabeled (PU) learning to address bias and ensure equitable outcomes in machine learning models trained on partially labeled data. Additionally, the study by (Hedderich et al., 2025) employed a human-centered framework, focusing on explicit linguistic cues and extracting token-level patterns that highlight systematic shifts, such as the use of gendered pronouns.
@@
More recently, there has been a shift towards addressing the implicit biases, which are not overtly expressed but encoded through subtle cues. Studies such as (Wen et al., 2025; Borah and Mihalcea, 2024; Kumar et al., 2024; Aldayel et al., 2024; Tan and Lee, 2025) analyze the presence of implicit biases in single-turn conversations, revealing that LLMs frequently fail to flag or respond adequately to covertly prejudiced language. Another study by (Rescala et al., 2024) used the 2019 argument dataset to examine the LLMs’ responses (single-turn) and their convincing attributes. A recent study by (Lake et al., 2025) analyzed the post-alignment distributional shift of LLM responses using open-ended QA datasets. The study finds that alignment reduces surface-level diversity while increasing the comprehensiveness of single responses. Thus, they define the stance as the response confirmation of the question-answer as “both”, “yes”, or “no”. Arora et al. Hofmann et al. (2024) frame the implicit racial bias in LLMs by prompting models with identity-linked names and contexts, revealing disparities in sentiment and response quality across demographic groups. The study by (Tan et al., 2025) explores model alignment through the analysis of implicit preferences as latent social values, which are inferred from community engagement patterns found in user-generated content. Ryan et al. (2024) examined the effect of aligning language models to specific preference sets and shows that the alignment of language models is not a One-Size-Fits-All. Multi-turn conversational stance dynamics have also been explored, as seen in Flek, Venkata Charan Chinni and Manish Gupta and Lucie et al. , where “dogmatism” is assessed through evolving stances. More precisely, the study tracks how users shift their stances across Reddit conversations and classifies their overall dogmatism based on these evolving stances.
@@
Framing Implicit Opinion Through Subtle Language. Upon examining the effect of implicit language, prior work has explored how subtle cues influence the interpretation of tasks, such as the interpretation of superlative comparisons (Pyatkin et al., 2025). A notable line of research investigates the general effects of linguistic subtlety, such as the use of superlatives or indirect references (Pyatkin et al., 2025; Liu et al., 2023). Another work extensively studied the identification of implicit hatespeech (Sap et al., 2020; ElSherief et al., 2021) or Sarcasm detection in dialogue using subtle cues (Ghosh et al., 2017). In these studies, implicitness is often assessed based on the surface representation, on whether the target group is explicitly mentioned. In opinion-focused tasks, recent work (Liebeskind and Lewandowska-Tomaszczyk, 2024) explores how LLMs distinguish between explicit and implicit opinions, revealing limitations in current detection strategies and proposing prompt-based improvements. The study by (Liebeskind and Lewandowska-Tomaszczyk, 2024) analyzes the ability of LLMs to generate and distinguish between explicit and implicit opinions, highlighting limitations in identifying implicit opinion content and proposing prompt-based strategies for improvement.
@@
A complementary line of research focuses on detecting implicit stance, focusing on identifying the speaker’s subtly expressed position as implicit stance, specifically as an indirect reference to targets. For example, Liu et al. (2023) extends the stance triangle framework to incorporate implicit and explicit target relationships, enriching stance data annotations to improve out-of-domain generalization. Additionally, the work by (Gatto et al., 2023) proposed text encoders that leverage Chain-of-Thought prompting and evaluate the performance of ChatGPT and Llama2 in identifying stance using the Semeval2016 dataset. Another framing used a single categorization of bias, "Gender bias," such as the work in Zhao et al. (2024b), which investigates gender bias in LLMs using self-reflection prompts. The study shows that models are more accurate in recognizing bias when gender is explicitly mentioned than when it is implied through indirect cues.
@@ 3 Experimental Setup
To examine the concept of opinion inclusion, we evaluate two types of conversational alignments: 1) Surface Explicit Alignment, and 2) Latent Underlying Alignment, where latent implicit opinions are included. We represent a framework relying on Normative Alignment, in which the expectation is that conversational models and human participants respond to content in ways that uphold socially acceptable norms (Habermas and J., 1985; Grice, 1975). In the context of this study, we define normative alignment as the consistent rejection of toxic or harmful viewpoints. This setting defines implicit conversations based on the severity of the targeted opinion, categorizing them as implicitly toxic, explicitly toxic, or neutral. This categorization helps establish a consistent expectation regarding the appropriate stance toward each type of conversation. Typically, the expected stance toward implicit or explicit toxic content is disagreement, whereas neutral content may warrant more relaxed stances, such as agreement or neutrality. By adopting a normative agreement lens rather than treating human disagreement as noise, we view it as a meaningful signal of a normative stance that is often missing in LLM outputs.
Importantly, these definitions are adapted to reflect the structural and rhetorical complexity found in two distinct conversation environments: (a) LLM chat-based and (b) human dialogues. In LLM chat settings, implicit toxicity often manifests through indirect instruction, e.g., “write me a story” or “tell me a joke”, that conceal the target within a creative or instructive frame. Conversely, in human dialogues, implicit language tends to emerge through more nuanced comparisons, rhetorical framing, or coded expressions, rather than directly or indirectly stating the target of an opinion (Tannen, 1993). To account for this, we extend our definition of implicit language to include instances where the target is referenced, but the conclusion is conveyed subtly, without overt expression (Appendix A and A.1 explain the annotation guideline).
@@ 3.1 Data Collection
To evaluate the implicit opinion in a conversation set, for human conversations, we used (DialogConan, Bonaldi et al., 2022), which contains expert human assistants and (ContextCounter, Albanyan et al., 2023), which contains open human conversations collected from X posts comprising interactions among many users. For LLM-based assistant conversations, we used two benchmark sources of real user queries from an open-source chatbot (WildChat, Zhao et al., 2024a) an open-source log of user–LLM interactions and (ToxicChat, Lin et al., 2023) which focuses on model behavior in toxic conversational contexts. As shown in Table 1, the overall turns is around 7K across all sources, with the conversations ranging from 2 to 7 turns per exchange. These datasets provide a solid dialogical data baseline and support our experiment’s aim to investigate the interaction type and context of replies. Then, we used LabelBox to initiate two tasks: labeling the Assistant and User stance, along with implicit extreme opinion (implicitly or explicitly toxic opinion). Details are provided in Appendix A.1.
@@ 3.2 Inclusive Implicit Learning models
We evaluate two learning paradigms to assess the model’s ability to internalize subtle opinion cues: 1) post-training using Instruct Tune on implicit conversations using decoder-only LLMs and 2) positive-unlabeled (PU) online learning using linear and shallow neural models trained on Sentence-BERT embeddings. In both setups, the training data includes varying proportions of implicit opinion examples, ranging from 10% to 100%, to evaluate scalability and robustness. Zero-shot and 0% implicit training settings are included as lower baselines. As the implicit opinions usually remain unlabeled or are harder to annotate. This case of scarcity of unlabeled examples has been extensively studied as a Positive-Unlabeled (PU) learning scenario (Jung and Wang, 2024a), with a focus on explicitly mentioning the target group. Instead, our study examines another angle of implicit and subtle reference to opinion. Thus, we formulate positive samples to include explicitly labeled stances, while unlabeled samples include texts with potential implicit stances (which might be Agree or Disagree). We formulate our task as a binary stance classification problem between Agree (positive class) and Disagree (negative class). Only these two stance categories are retained during preprocessing. In (PU) training, for each assistant response, we concatenate the user and assistant messages (user [SEP] assistant) and represent them using dense semantic embeddings from a Sentence-BERT model (all-MiniLM-L6-v2). As for LLMs (Llama3 and Mistral), we used an instruction tuning prompt that includes the context of user implicit opinion (Appendix C).
@@
We adopt principles from positive-unlabeled learning (Jung and Wang, 2024a) to handle imbalance and fairness settings between implicit and explicit contextual opinion expressions. Each example is tagged with a sensitive attribute based on whether the user message expresses an implicit (represented as 0) or explicit (represented as 1) opinion. These group indicators are used to handel fairness constraints in PU, ensuring that models maintain comparable false positive rates (FPR) across both implicit and explicit opinions.

### v2

@@ Abstract
Shaping inclusive representations that embrace diversity and ensure fair participation and reflections of values is at the core of many conversation-based models. However, many existing methods rely on surface inclusion using mention of user demographics or behavioral attributes of social groups. Such methods overlook the nuanced, implicit expression of opinion embedded in conversations. Furthermore, the over-reliance on overt cues can exacerbate misalignment and reinforce harmful or stereotypical representations in model outputs. Thus, we took a step back and recognized that equitable inclusion needs to account for the implicit expression of opinion and use the stance of responses to validate the normative alignment. This study aims to evaluate how opinions are represented in NLP or computational models by introducing an alignment evaluation framework that foregrounds implicit, often overlooked conversations and evaluates the normative social views and discourse. Our approach models the stance of responses as a proxy for the underlying opinion, enabling a considerate and reflective representation of diverse social viewpoints. We evaluate the framework using both (i) positive-unlabeled (PU) online learning with base classifiers, and (ii) instruction-tuned language models to assess post-training alignment. Through this, we provide a principled and structured lens on how implicit opinions are (mis)represented and offer a pathway toward more inclusive model behavior.
@@ 1 Introduction
Recent studies have begun to examine the implicit bias behavior of models, particularly in scenarios where bias is conveyed through covert or subtle linguistic cues (Hofmann2024-ml; aldayel-etal-2024-covert). Given that social norms are situational and bias remains contextual, this urges a need for a scheme that places these considerations at the core of the process (Wen2025-wc). Thus, we take a step back to evaluate how implicit opinions are contextually expressed and interpreted within conversational settings. This aspect is based on the Implicit Attitude Theory, which indicates that individuals hold attitudes that may not be explicitly expressed but are reflected in implicit ways (greenwald1995implicit-f4c). Following Grice’s Cooperative Principle (grice1975logic), which explains how meaning is often conveyed through implicature and indirectness, we consider how speakers may express minority or dissenting viewpoints implicitly or indirectly, in ways that adhere to social expectations while avoiding overt conflict. On the light of these theoretical foundations, the EMBRACE framework (Engaging Multiple Beliefs, Reflections, and Contexts Equitably) emphasizes the importance of surfacing and incorporating implicit viewpoints during model training and evaluation. More practically, the inclusion of implicit conversational turns enhances stance norm alignment by allowing models to learn pragmatic inference patterns rather than relying solely on surface-level agreement indicators. This framework can help explain the tendency of LLMs to inadequately represent diverse perspectives and opinions, as their training data often underrepresents implicit or indirect expressions of opinion.
@@
Many previous methods on pluralistic opinions (Feng2024-ny; Sorensen2024-zj) have focused on superficial characteristics, without a careful distinction between related yet distinct concepts, opinion and stance. Opinion refers to individual’s subjective belief or attitude about a topic or entity. It often reflects a speaker’s evaluation, which may be explicit or implied in language (Oskamp2005-li). While Stance, in contrast, refers to the speaker’s expressed position or orientation toward a specific proposition or opinion. Stance is often shown through agreement, disagreement, or neutrality in response to another utterance (Bois2007-ps; ALDayel2021-fj). Therefore, in conversation, a stance is observable alignment that may reflect an opinion, but it can also be situational. In this way, opinions can inform stances, but they remain latent unless made explicit through discourse.
@@
To this end, we evaluate how the implicit opinion affects the follow-up stance in this work. We present a framework to assess the impact of implicit opinion in discourse. We examine how stance and certainty cues manifest differently in implicit versus explicit opinionated conversations to uncover subtle patterns of opinion expression. First, we establish the framework to validate normative alignment, in which a unified expectation guides appropriate responses for equitable inclusion. This expectation stems from normative discourse principles (habermas2015theory; grice1975logic), where toxic language (e.g., hate, dehumanization, or extreme ideological views) is not treated neutrally but is instead met with opposition. By aligning stance judgments with this expectation, we can measure whether models reinforce or resist harmful views, especially when they are expressed implicitly. Then, we highlight key turning points in multi-turn dialogues where stance certainty changes, providing insights into how opinions evolve throughout the conversation. Finally, we show that incorporating implicit turns into computational models affects stance classification performance, illustrating how such inclusion can either amplify or mitigate the expression and identification of opinions.
@@ 2 Related work
Opinion and Bias Representation. Implicit opinion bias has been defined as the use of subtle language, including hedging, implicature, and abstraction, which can preserve or amplify social stereotypes even in the absence of explicit prejudice (Maass1999-ay; Tannen1993-nf). Most previous work on opinion and bias has focused on direct, explicit social biases, such as gender disparities in word embeddings (Cheng2022-tn) or demographic biases in LLMs (Hedderich2025-se). Several studies have also examined the racial aspect of bias (Hofmann2024-ml; Sun2025-vu), often operationalized through identity-linked prompts or response disparities on tone or sentiment polarity. For instance, the study by (jung2024fairness-aware-5eb) developed fairness-aware methods for online Positive-Unlabeled (PU) learning to address bias and ensure equitable outcomes in machine learning models trained on partially labeled data. Additionally, the study by (Hedderich2025-se) employed a human-centered framework, focusing on explicit linguistic cues and extracting token-level patterns that highlight systematic shifts, such as the use of gendered pronouns.
@@
More recently, there has been a shift towards addressing the implicit biases, which are not overtly expressed but encoded through subtle cues. Studies such as (Wen2025-wc; borah2024towards-ea8; Kumar2024-ru; aldayel-etal-2024-covert; Tan2025-dp) analyze the presence of implicit biases in single-turn conversations, revealing that LLMs frequently fail to flag or respond adequately to covertly prejudiced language. Another study by (rescala2024can-ecd) used the 2019 argument dataset to examine the LLMs’ responses (single-turn) and their convincing attributes. A recent study by (lake2025from-f40) analyzed the post-alignment distributional shift of LLM responses using open-ended QA datasets. The study finds that alignment reduces surface-level diversity while increasing the comprehensiveness of single responses. Thus, they define the stance as the response confirmation of the question-answer as “both”, “yes”, or “no”. Arora et al. Hofmann2024-ml frame the implicit racial bias in LLMs by prompting models with identity-linked names and contexts, revealing disparities in sentiment and response quality across demographic groups. The study by (Tan2025-xb) explores model alignment through the analysis of implicit preferences as latent social values, which are inferred from community engagement patterns found in user-generated content. ryan2024unintended-6e2 examined the effect of aligning language models to specific preference sets and shows that the alignment of language models is not a One-Size-Fits-All. Multi-turn conversational stance dynamics have also been explored, as seen in fleknoyearusdc-236, where “dogmatism” is assessed through evolving stances. More precisely, the study tracks how users shift their stances across Reddit conversations and classifies their overall dogmatism based on these evolving stances.
@@
Framing Implicit Opinion Through Subtle Language. Upon examining the effect of implicit language, prior work has explored how subtle cues influence the interpretation of tasks, such as the interpretation of superlative comparisons (pyatkin-etal-2025-superlatives). A notable line of research investigates the general effects of linguistic subtlety, such as the use of superlatives or indirect references (pyatkin-etal-2025-superlatives; Liu2023-ws). Another work extensively studied the identification of implicit hatespeech (Sap2020-bp; ElSherief2021-br) or Sarcasm detection in dialogue using subtle cues (Ghosh2017-hh). In these studies, implicitness is often assessed based on the surface representation, on whether the target group is explicitly mentioned. In opinion-focused tasks, recent work (liebeskind2024navigating-937) explores how LLMs distinguish between explicit and implicit opinions, revealing limitations in current detection strategies and proposing prompt-based improvements. The study by (liebeskind2024navigating-937) analyzes the ability of LLMs to generate and distinguish between explicit and implicit opinions, highlighting limitations in identifying implicit opinion content and proposing prompt-based strategies for improvement.
@@
A complementary line of research focuses on detecting implicit stance, focusing on identifying the speaker’s subtly expressed position as implicit stance, specifically as an indirect reference to targets. For example, Liu2023-ws extends the stance triangle framework to incorporate implicit and explicit target relationships, enriching stance data annotations to improve out-of-domain generalization. Additionally, the work by (Gatto2023-dt) proposed text encoders that leverage Chain-of-Thought prompting and evaluate the performance of ChatGPT and Llama2 in identifying stance using the Semeval2016 dataset. Another framing used a single categorization of bias, "Gender bias," such as the work in Zhao2024-mv, which investigates gender bias in LLMs using self-reflection prompts. The study shows that models are more accurate in recognizing bias when gender is explicitly mentioned than when it is implied through indirect cues.
@@ 3 Experimental Setup
To examine the concept of opinion inclusion, we evaluate two types of conversational alignments: 1) Surface Explicit Alignment, and 2) Latent Underlying Alignment, where latent implicit opinions are included. We represent a framework relying on Normative Alignment, in which the expectation is that conversational models and human participants respond to content in ways that uphold socially acceptable norms (habermas2015theory; grice1975logic). In the context of this study, we define normative alignment as the consistent rejection of toxic or harmful viewpoints. This setting defines implicit conversations based on the severity of the targeted opinion, categorizing them as implicitly toxic, explicitly toxic, or neutral. This categorization helps establish a consistent expectation regarding the appropriate stance toward each type of conversation. Typically, the expected stance toward implicit or explicit toxic content is disagreement, whereas neutral content may warrant more relaxed stances, such as agreement or neutrality. By adopting a normative agreement lens rather than treating human disagreement as noise, we view it as a meaningful signal of a normative stance that is often missing in LLM outputs.
Importantly, these definitions are adapted to reflect the structural and rhetorical complexity found in two distinct conversation environments: (a) LLM chat-based and (b) human dialogues. In LLM chat settings, implicit toxicity often manifests through indirect instruction, e.g., “write me a story” or “tell me a joke”, that conceal the target within a creative or instructive frame. Conversely, in human dialogues, implicit language tends to emerge through more nuanced comparisons, rhetorical framing, or coded expressions, rather than directly or indirectly stating the target of an opinion (Tannen1993-nf). To account for this, we extend our definition of implicit language to include instances where the target is referenced, but the conclusion is conveyed subtly, without overt expression (Appendix A and A.1 explain the annotation guideline).
@@ 3.1 Data Collection
To evaluate the implicit opinion in a conversation set, for human conversations, we used (DialogConan, bonaldi2022human-machine-cf8), which contains expert human assistants and (ContextCounter, albanyan2023not-b79), which contains open human conversations collected from X posts comprising interactions among many users. For LLM-based assistant conversations, we used two benchmark sources of real user queries from an open-source chatbot (WildChat, Zhao2024-wz) an open-source log of user–LLM interactions and (ToxicChat, Lin2023-zt) which focuses on model behavior in toxic conversational contexts. As shown in Table 1, the total number of turns is approximately 7K across all sources, with the conversations ranging from 2 to 7 turns per exchange. These datasets provide a solid dialogical data baseline and support our experiment’s aim to investigate the interaction type and context of replies. Then, we used LabelBox to initiate two tasks: labeling the Assistant and User stance, along with implicit extreme opinion (implicitly or explicitly toxic opinion). Annotators were trained on pilot examples using Labelbox interface, with consensus labeling of stance (Agree/Disagree/Neutral), opinion type (Implicit vs Explicit Toxicity), and Certainty (Certain, Uncertain, Refuse to Engage, None). The full annotation guidelines and anonymized examples are described in Appendix A.1, and the corresponding dataset is publicly available on OSF.111https://osf.io/2azn5/files?view_only=bd8f756bb2e849a1b5102953cf33a775.
@@ 3.2 Inclusive Implicit Learning models
We evaluate two learning paradigms to assess the model’s ability to internalize subtle opinion cues: 1) post-training using Instruct Tune on implicit conversations using decoder-only LLMs and 2) positive-unlabeled (PU) online learning using linear and shallow neural models trained on Sentence-BERT embeddings. In both setups, the training data includes varying proportions of implicit opinion examples, ranging from 10% to 100%, to evaluate scalability and robustness. Zero-shot and 0% implicit training settings are included as lower baselines. As the implicit opinions usually remain unlabeled or are harder to annotate. This case of scarcity of unlabeled examples has been extensively studied as a Positive-Unlabeled (PU) learning scenario (jung2024fairness-aware-5eb), with a focus on explicitly mentioning the target group. Instead, our study examines another angle of implicit and subtle reference to opinion. Thus, we formulate positive samples to include explicitly labeled stances, while unlabeled samples include texts with potential implicit stances (which might be Agree or Disagree). We formulate our task as a binary stance classification problem between Agree (positive class) and Disagree (negative class). Only these two stance categories are retained during preprocessing. In (PU) training, for each assistant response, we concatenate the user and assistant messages (user [SEP] assistant) and represent them using dense semantic embeddings from a Sentence-BERT model (all-MiniLM-L6-v2). As for LLMs (Llama3 and Mistral), we used an instruction tuning prompt that includes the context of the user’s implicit opinion. For reproducibility, we used Mistral-7B-Instruct-v0.3 and LLaMA-3-8B-Instruct models, each fine-tuned with LoRA adapters on implicit conversation data. Complete hyperparameters and model parameter sizes are summarized in Appendix C (Tables 7–8).
@@
We adopt principles from positive-unlabeled learning (jung2024fairness-aware-5eb) to handle imbalance and fairness settings between implicit and explicit contextual opinion expressions. Each example is tagged with a sensitive attribute based on whether the user message expresses an implicit (represented as 0) or explicit (represented as 1) opinion. These group indicators are used to handle fairness constraints in PU, ensuring that models maintain comparable false positive rates (FPR) across both implicit and explicit opinions.